OpenStudy (anonymous):

what do you do in this equation? (y+10)(y+5)=6

OpenStudy (anonymous):

first multiply if u need to solve or simplify

OpenStudy (anonymous):

foil method?

OpenStudy (anonymous):

y^2+15y +44 =0 or y^2+11y+4y+44=0 or(y+4)(y+11)=0 hence y=-4 or y=-11

OpenStudy (anonymous):

so you took away six??

OpenStudy (anonymous):

you looked for a perfect square?

OpenStudy (anonymous):

we simplified (y+10)(y+5) which gives y^2 + 15y + 50 then subtract the 6 and then factor so we can replace 15y with 11y +4y and factor again (y+11) (y+4)=0 y=-11 or y=-4

OpenStudy (anonymous):

ok thank you because I couldnt get this